### v1

Abstract
A central problem in the deployment of quantum technologies is the realization of robust architectures for quantum interconnects. We propose to engineer interconnects in semiconductors and insulators by patterning spin qubits at dislocations, thus forming quasi one-dimensional lines of entangled point defects. To gain insight into the feasibility and control of dislocation-driven interconnects, we investigate the optical cycle and coherence properties of nitrogen-vacancy (NV) centers in diamond, in proximity of dislocations, using a combination of advanced first-principles calculations. We show that one can engineer spin defects with properties similar to those of their bulk counterparts, including charge stability and a favorable optical cycle, and that NV centers close to dislocations have much improved coherence properties. Finally, we predict optically detected magnetic resonance spectra that may facilitate the experimental identification of specific defect configurations. Our results provide a theoretical foundation for the engineering of one-dimensional arrays of spin defects in the solid state.
1 Main
One of the key components of quantum technologies is the interconnect – an architecture or physical process that enables the transfer of information between disparate and varied quantum nodes [1, 2, 3, 4]. In the solid state, controlling and scaling defect-based qubits into coherent, interconnected arrays remain major obstacles, limiting the application of spin defects for quantum technologies [5, 6, 7, 8]. Here we address this challenge and propose to create quantum interconnects by using the quasi one-dimensional (1D) atomic arrangements present at dislocations to pattern and eventually entangle spin qubits in semiconductors and insulators.
Dislocations are topological, line-defects in crystals [9]. At their narrow core, atoms adopt complex arrangements that differ from those in the rest of the solid. The unique bonding environments of dislocation cores can give rise to emergent electronic and magnetic properties [10, 11, 12, 13]; in addition, dislocations can attract vacancies and substitutional impurities [14, 15, 16], due to their long-range stress/strain field, offering an opportunity to engineer unique structures of interacting line and point defects.
Recently, some preliminary theoretical studies explored the prospect of assembling arrays of spin defects [15, 16]. Close to the core of dislocations, the formation energies of nitrogen vacancy centers (NV) in diamond [15] and divacancies in silicon carbide [16] were found to be much lower relative to the bulk (up to 3-4 eV lower). Interestingly, Ref. 15 identified an NV configuration at the 30∘ dislocation core which exhibits stability properties similar to the corresponding bulk NV. These results suggest that spin defects may be patterned to form interconnects of qubits, exploiting the thermodynamic driving force provided by dislocations.
However, the viability of spin defects as qubits close to dislocations, and specifically their optical cycle and coherence properties, have never been investigated, and the impact of line defects on their functionality remains unknown. In fact, it is only recently that advanced computational methods to study the full optical cycles of NV centers, including inter-system crossing (ISC) rates, became available [17] and that calculations beyond mean-field, density functional theory, could be applied to systems complex enough to describe, at the same time, dislocations and point defects [18]. To provide a theoretical foundation for the engineering of 1D arrays of spin defects in the solid state, we present a detailed investigation of the optical cycle of NV centers close to dislocations as a representative system, using a combination of recently developed first-principles methods and large-scale simulations. Our investigation integrates electronic structure and coherence time calculations, uncovering novel dynamical properties of spin defects close to dislocations, in addition to providing predictions on how to leverage the interaction of spin defects with line dislocations to enhance their properties.
Through a high-throughput search of stable centers near specific dislocations, we identify triplet configurations stable in the negative (i.e., ) charge state. We then carry out an in-depth investigation of the optical cycle of representative NVs. We show that despite the variability of properties depending on the geometrical configuration, one can engineer spin defects near dislocations that not only retain the essential characteristics required for qubit functionality, but also exhibit longer coherence times compared to those in pristine diamond.
NV centers close to dislocation cores
We focus on the 30∘ and 90∘ glide partial dislocations, which arise from the energetically favorable dissociation of the 60∘ glide perfect dislocations [19, 20]. Specifically, we consider the double-period reconstruction of the core geometry [13], which reduces the energy of the system by minimizing the number of dangling bonds (Fig. S1). We carry out first principles calculations with a combination of advanced techniques, ranging from density functional theory (DFT) and time-dependent (TD) DFT with hybrid functionals and analytical forces [21] to quantum defect embedding [22, 23] and cluster-expansion for coherence times calculations [24, 25]. We start by describing the stability of NV centers, and then we discuss their electronic properties and optical cycle.
Fig. 1 (a-d) shows the core structures of two dislocations and the atomic sites considered to create NV defects near the core region, and (e-f) the formation energy of all NV configurations analyzed here [see Supplementary Information (SI) for details of the enumeration of NV configurations]. In total, we screened 90 (112) NV arrangements near the 30∘ (90∘) dislocation core considering all atomic sites colored in orange, red, and blue in Fig. 1 (c) [(d)]; we optimized their geometries in large supercells (Fig. S1) using DFT with the Perdew-Burke-Ernzerhof (PBE) functional [26], in either the singlet or triplet spin state, obtaining their relative formation energy (). Consistent with previous reports [15], we find that most of the NV centers have a lower formation energy () than in the bulk [Figs. 1 (e-f), S2, and S3]. A substantial fraction are also stable in the triplet spin state. Among the 90 NVs near the 30∘ dislocation core, 62 (69%) favor the triplet spin state over the singlet, with 46 (51%) exhibiting a triplet state that is at least 0.1 eV lower in energy. Similarly, for the 112 NVs near the 90∘ dislocation core, 76 (68%) favor the triplet spin state, with 57 (51%) exhibiting a triplet state that is at least 0.1 eV lower in energy. Since not all centers considered here are stable in the desired triplet spin state, in a typical sample, some of the defects attracted to dislocations will be dark.
Among the screened configurations with low formation energies, we then selected a few representative NV defects (Fig. S4) that are stable in a negative charge state and a triplet spin state, at least 0.1 eV lower in energy than the singlet state (to avoid possible temperature-induced transitions); their properties were examined in detail as discussed below.
All selected representative NVs are metastable compared to the most stable one, i.e., the G-9-G11 (G-6-G9) near the 30∘ (90∘) dislocation core; however, we expect that metastable configurations may be formed as a result of kinetic trapping. To understand the charge stability of these NVs, we computed their formation energy as a function of the system Fermi level (), and obtained the charge transition levels (CTL). We used the dielectric-dependent hybrid (DDH) functional [27] and considered PBE optimized geometries. We find that all representative defects have lower than in bulk diamond, for all possible values within the gap (Fig. S5). Despite the variations of CTL positions, all representative defects can be stable in charge states from +1 to depending on the chosen value of . Importantly, they can be stabilized in the charge state for a wide range of . An effective way to tune the Fermi level could be, for example, through the control of N dopants. We note that the dislocations studied in this work do not interfere with the tuning and controlling of (Fig. S5) as they do not introduce deep states in the band gap of the solid.
Fig. 2 presents a comparison of the single-particle states of the bulk NV and the G9-G8 (G4-G9) NV at the 30∘ (90∘) dislocation. Results for other representative defects are presented in Figs. S6 and S7. As expected, the symmetry of the defects is lowered close to dislocations, and thus the defect states, degenerate in bulk diamond, are split into two distinct levels [Fig. 2 (b) and (c)]. Not surprisingly, this splitting is observed for all representative NVs, and its magnitude, as well as the position of the defect levels within the gap, vary depending on the configuration.
Nevertheless, several important properties for the functionality of NV centers as qubits are shared by all configurations: the highest occupied () and the lowest empty defect level () in the spin-down channel are well separated from the electronic states of the host (Figs. 2, S6, and S7). Further, the energy spacing between the and levels is sufficiently large (2 eV) to prevent intra-defect thermal excitations near room temperature. Finally, the strain fields and distinct bonding geometries of dislocations give rise to electronic states within the gap of bulk diamond that are delocalized along the dislocation core and close in energy to the band edges of the solid. We find that these states are well separated from the defect levels (Figs. 2, S6, and S7), as also shown by the analysis of localization factors reported in Table S1.
Excited-state properties of NV centers at dislocations
We now turn to describe the excited state properties of the representative NVs introduced above, computed using TDDFT with the DDH functional. These calculations employed recently developed methods and optimized GPU-accelerated codes [18, 21]. In particular, we computed and analyzed the intra-defect vertical excitation energies (VEEs) between many-body states. Fig. 3 shows a comparison of the many-body state diagrams of the bulk NV and NV centers near dislocations; the diagrams for the other representative NVs are presented in Figs. S8 and S9. We observe two triplet excited states [(2) and (3)] originating from spin-conserving excitations within the spin-down channel, where an electron is excited from to one of the split and levels (Fig. 2). The spin-flip excitations from and to and lead to transitions from (1) to (1), (2) and (3) states.
Similar to single-particle states, the many-body state diagrams of the NVs near the dislocation core vary depending on the configuration. However, for all representative spin defects, except for the G-9-G-10 near the 30∘ dislocation, the (1) (2) transition is the lowest spin-conserving vertical excitation occurring in the spin-down channel, and the corresponding VEEs are smaller than that of defect-to-band transitions (Table S2), similar to what was found for the bulk NV. Hence, our calculations show that a key requisite to prepare and measure the qubit state is satisfied for NV centers close to dislocations: the intra-defect optical transitions do not interfere with the electronic states of the host. Further, we observe three singlet excited states [(1), (2) and (3)] positioned between the two triplet states (1) and (2) (Figs. 3, S8 and S9), indicating that an optical initialization cycle could in principle be realized that resembles that of the bulk NV. However, in some cases, e.g., the G3-G8 configuration near the core of 30∘ dislocation (Fig. S8), the singlet excited state (3) is higher in energy than the lowest triplet excited state (2), and thus unlikely to be part of the non-spin-conserving decay path of the (2) state. We also find that the VEEs from the ground triplet to the lowest singlet excited state for NVs near the dislocation core region are all smaller than in the bulk, suggesting an increased stability of singlet spin states near the core. Our calculations further show that the VEEs for the (1) (1) transition are generally larger than the thermal energy at room temperature (0.025 eV), thus preventing thermal excitations; however, there are exceptions with a small VEE, e.g., the G10-G9 NV near the core of 90∘ dislocation, and these configurations are not expected to be useful qubits at least near room temperature.
We also computed the zero-phonon line (ZPL) energy for the spin defects in the bulk and at dislocations. Specifically, we computed the ZPL for the transition from the first triplet excited state to the ground state [ or (2) (1); Figs. 3, S8, and S9]. Experimentally, ZPL values can be obtained from photoluminescence (PL) spectra. Hence, our results not only help characterize NVs near the core region, but also assist in their experimental identification. Fig. S10 summarizes the computed ZPL values and the Debye-Waller factors (DWF). Depending on the configuration and its proximity to the dislocation core, the computed ZPL can vary substantially (from 1.06 to 2.49 eV), compared to that (2.07 eV) in bulk diamond. Similar to the ZPL, NVs at dislocations exhibit varied DWFs. The largest (smallest) DWF obtained for the representative NVs is 10.17% (0.55%), compared to that (4.89%) in bulk diamond, showing that all these defects may be observed by experiments with a detectable ZPL signal.
Inter-system crossing rates and optical cycle
To understand the optical spin initialization and readout processes of NVs near the dislocation core, we computed the ISC rates of three prototypical configurations; our results are shown in Fig. 3. As mentioned in the introduction, we employed here methodologies and codes that only recently became available to study ISC processes with advanced first-principles techniques [17]. We considered the experimental values for the bulk NV as reference [28]; we then computed the variation of the spin-orbit coupling (SOC) matrix elements between the triplet and singlet states using the many-body wavefunctions obtained from the quantum defect embedding theory (QDET) [22, 23], and the vibrational overlap functions for the bulk NV, at the respective estimated zero-phonon transition energies [17] (see SI). The symmetry lowering of the spin defects close to dislocations, from to , lifts the zero-field degeneracy of the states, which hybridize to form the non-degenerate levels . Fig. 3 shows the many-body electronic states diagrams for bulk NV and NVs near the core of dislocations. For simplicity, here we only consider the ISC transitions between and (denoted as upper ISC), and between and (lower ISC). The ISC transition rates from to , and from to are assumed to be small compared to the non-radiative transition from to and from to , given the small energy gap between these levels.
We recall that for the bulk NV, the upper ISC rates show a preferential selectivity between the sub-level of and , which is key to its optical spin initialization and readout processes, while the lower ISC rates are much smaller and less selective. We find that the ISC rates of three NVs near the dislocation core analyzed here show quantitatively different behaviors, as illustrated in Fig. 3 (b-d). For the G9-G8 configuration near the core of 30∘ dislocation [Fig. 3 (b)], the upper ISC rate is much smaller than in bulk NV, due to the increased energy gap between and ; the selectivity for the ISC transitions from sub-levels to is similar to bulk NV; the lower ISC rates are faster than bulk NV, due to the decreased energy gap between and , and the selectivity is higher for the spin sub-level. For the G4-G9 configuration near the core of 90∘ dislocation [Fig. 3 (c)], the upper ISC rates are comparable in magnitude to those in bulk NV, and they show an ISC selectivity from the sub-level of to ; the lower ISC rate is also selective from the to the sub-level of and faster than in bulk NV. For the G+7-G+3 configuration near the core of 30∘ dislocation [Fig. 3 (d)], the upper ISC rates are comparable in both magnitude and selectivity to bulk NV; the lower ISC rates are much faster, given the decreased energy gap between and and increased SOC matrix elements, due to the breaking of the symmetry. Based on the zero-phonon energy gaps and vibrational overlap functions of the 18 representative NVs (Fig. S11), we estimate four of them (22%) to exhibit desirable ISC rates.
Importantly, the essential properties required to enable an optical cycle are maintained for some NVs close to the core, as shown by state population simulations in Fig. 4. We considered a seven-state model (Fig. S12): , , and (, , and for the bulk NV) spin sub-levels of the triplet ground state (GS), triplet excited state (ES), and one singlet state (SS); the latter is chosen to encompass the , , and singlets ( and singlets for the bulk NV) considering the relatively fast non-radiative transition rates between the various singlet states [17]. We performed three types of simulations for the optical spin initialization, the continuous wave optically detected magnetic resonance (cw-ODMR), and the PL spin readout (see SI for details).
From Fig. 4 (a), we see that the bulk NV can be initialized into the spin sub-level up to about 70%, due to the selectivity of its upper ISC. Its ODMR contrast can reach about 18% under the simulation conditions, showing its applicability in quantum sensing. Its PL spin readout shows a distinct difference for initialization into the or sub-levels, thus allowing for the readout of the population on different spin sub-levels using time-resolved PL measurements. These results are in agreement with what was previously reported for NVs in bulk diamond [29, 30, 31].
For the G9-G8 configuration near the core of 30∘ dislocation [Fig. 4 (b)], the initialization takes 100 times longer to reach the steady state than in bulk NV, due to the much smaller upper ISC rates, and the ODMR contrasts are 40 times smaller; its time-resolved PL displays negligible differences between different initialization conditions, implying that this NV may not be useful for quantum applications based on optical control. The G4-G9 configuration near the core of 90∘ dislocation [Fig. 4 (c)] can be initialized to the sub-level up to about 75% because of the much faster lower ISC to the GS sub-level; its ODMR has a positive contrast, given the reversed selectivity of the upper ISC, compared to that of the bulk NV; its time-resolved PL exhibits a noticeable difference between initialization into the or sub-level. The G+7-G+3 configuration near the core of 30∘ dislocation [Fig. 4 (d)] exhibits an initialization time similar to the bulk NV; it has a small ODMR contrast between the and sub-levels because of the comparable triplet ES to SS and triplet ES to SS ISC rates; its time-resolved PL displays a noticeable difference between the and sub-levels and can be potentially used for PL spin readout.
The state population simulations show that there exist NV configurations close to dislocations that can be properly initialized [Fig. 4 (c) and (d)] and used in an optical cycle similar to that of the bulk. In addition, our results provide a means to experimentally identify configurations with specific characteristics close to dislocations.
Coherence times
Finally, we discuss coherence times of representative NV defects close to dislocations, and we start by evaluating their axial () and transverse () components of the ground-state zero-field splitting (ZFS) tensor (Fig. S13). The highest (lowest) value is found to be 3839.0 (1796.6) MHz, deviating by up to 44% from that of the bulk NV (3230.8 MHz). The lowering of the defect symmetry leads to large nonzero values for the defects close to dislocations, with the highest predicted value reaching 813.3 MHz. Hence, the zero‐field degeneracy of the states is lifted to form non‐degenerate “clock” levels with no first‐order magnetic moment. As a result, NV centers at dislocations exhibit first‐order‐insensitive transitions in weak magnetic fields (), known as “clock transitions”. We computed the Hahn‐echo coherence time for magnetic fields up to mT [Figs. 5 (a), S14, and S15] and observe that decreases with increasing , reflecting enhanced decoherence from the field‐dependent dynamics of the 13C spin bath as one moves away from the zero‐field clock transition. This behavior is notably different from that of bulk NV (whose component of the ZFS is zero), which monotonically increases with and saturates at 1 ms [Fig 5 (a)] [32].
The representative NV centers near dislocations exhibit longer coherence times than the bulk NV, at 0 mT up to 1 mT depending on the ZFS parameters (Figs. S14 and S15). The observed enhancement of at is determined by both the and components of the ZFS tensor. As illustrated in Fig. 5 (b), an increase in and leads to a significant enhancement in , reaching up to an order of magnitude improvement compared to bulk NV. A similar enhancement is also observed in the free induction decay time (Fig. S16). To further investigate the protection of spin coherence from environmental noise, we simulated NV coherence as a function of total evolution time under periodic Carr-Purcell-Meiboom-Gill (CPMG) pulse sequences at on representative NVs near dislocations. As shown in Fig. 5 (c), the coherence time of the - NV at the dislocation increases from 3.7 ms to 100 ms with -pulses, and the traces in Fig. 5 (d) show that raising suppresses decoherence and steepens the decay. We note that the reported values were computed in the long- limit, where longitudinal relaxation does not affect NV coherence [33, 34]. Our results demonstrate that the ZFS tensor of the NV can be engineered via dislocations, significantly enhancing coherence time and providing optimal working points for NV qubits.
Discussion
We investigated the feasibility of using NV centers as spin qubits close to dislocations in diamond. We conducted a high-throughput screening of all possible NV configurations near the core regions of two types of dislocations, and showed that multiple configurations near the core can be stabilized in a triplet spin state and charge state. We then selected representative NVs and thoroughly examined their single-particle defect levels, optical cycle, and coherence times. In spite of differences in these properties relative to the bulk NVs, we found representative spin defects in the core region that meet the requisites to be useful qubits. Importantly, the optical cycle for qubit initialization and readout is feasible for several NVs. Moreover, certain centers can outperform the bulk NV, exhibiting a larger DWF and improved coherence time.
We emphasize that the optical cycle of spin defects close to dislocations has never been investigated before; we could address this problem here thanks to the integration of a comprehensive suite of accurate and scalable simulation techniques that have only recently been developed [17, 18, 21, 23], enabling the investigation of the interplay between extended and point defects. Importantly, the workflow established here can be readily extended to explore other defects and hosts. We also note that given the large system size required to account for both point and extended defects (we used a 1728-atom supercell of diamond), our massively parallel GPU-accelerated implementation of TDDFT and QDET in the WEST code [21, 23, 35, 36] turned out to be critical to obtain accurate results, as was our recently developed workflow [17] to predict ISC rates.
Based on the promising properties found in our study, we propose that NVs close to dislocations may be utilized to pattern and couple 1D arrays of spin defects; this engineering strategy to form quantum interconnects may advance the development of solid-state quantum technologies. Such a strategy is not limited to the system of NV qubits in diamond and could be utilized for other hosts such as SiC and other geometries, including semiconductor/oxide interfaces [37, 38].
However, clearly some challenges remain to be addressed to fulfill the potential of the engineering strategy proposed here. Only a fraction of spin defects close to dislocations display the desired quantum properties. Hence, a controlled synthesis and manipulation of spin defects with favorable functionalities need to be developed. Part of the synthetic strategy may include the controlled introduction and patterning of specific dislocations into the sample, as well as the use of impurities (e.g., N) to engineer the desired Fermi level values.
Methods
Using the 1728-atom supercells, we carried out a high-throughput search of configurations of NV centers close to dislocations using DFT as implemented in the Quantum ESPRESSO code [39]. We then computed their electronic and coherence properties, and compared them with those of NVs in bulk diamond. Using DFT, we computed the structural properties of spin defects with the PBE [26] functional, and formation energies in varied charge states, charge transition levels, and single-particle energies with the DDH functional [27]. We computed many-body energy levels using (spin-flip) TDDFT with the DDH functional, as implemented in the WEST code [21]. ISC rates were obtained with a newly developed framework [17] that combines calculations of many-body states with QDET [22, 23], fully relativistic calculations of SOC, and calculations of overlap functions using first-principles phonons [40, 41]. We used a seven-state model for state population simulations of ODMR spectra. Coherence times were obtained with the generalized cluster expansion techniques (gCCE) [24, 25] and the PyCCE code [42], and calculations of ZFS parameters were carried out with the PyZFS code [43]. Details are provided in the SI.
Supplementary Information Supplementary Text, Figs. S1–S16, Tables S1–S4 and References.
Acknowledgements We thank F. Joseph Heremans, Alex High, Benchen Huang, and Marquis M. McMillan for useful discussions. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-05CH11231. This research also used resources of the Argonne Leadership Computing Facility at Argonne National Laboratory, a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-06CH11357. This work was supported by the AFOSR Grant No. FA9550-23-1-0330. Software development (WEST, PyCCE, and PyZFS) in this work was supported by MICCoM, as part of the Computational Materials Sciences Program funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences, and Engineering Division through Argonne National Laboratory.